OpenStudy (anonymous):

Which of the following is equivalent to ab = 7?

OpenStudy (anonymous):

a=7/b

OpenStudy (anonymous):

b=7/a

OpenStudy (anonymous):

logb^7=a???

OpenStudy (anonymous):

hmm did u mean a^b=7?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

b = loga7 i think.. im rusty on logs.. @angela210793 take a look

OpenStudy (anonymous):

yes that is correct. confirmed